Three times the square of a number minus 6 is equal to 14.determine the value of the number
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52824)

.
This can be read in two ways:

1)  3x^2 - 6 = 14,     and/or


2)  3*(x-6)^2 = 14.

And the post does not give me the opportunity to choose between these two.



Answer by Boreal(15235)

number is x, 3 times it squared is 3x^2, 3 times it squared minus 6 is 3x^2-6
3x^2-6=14
3x^2=20
x^2=20/3
x=sqrt(60)/3, after rationalizing.
square x and it is 60/9
multiply it by 3 and get 60/3=20
20-6=14
numerically, x=2.58
=====================